OpenStudy (anonymous):

identify the x and y intercepts for the equation 4x – 3y = -12.

OpenStudy (lalaly):

for the x intercepts, set y=0 4x-3(0)=-12 4x=-12 x=-3 (-3,0) for the y intercepts set x=0 4(0)-3y=-12 -3y=-12 y=4 (0,4)